# subfigure0

### v1

@@ 2.3 Deformations
The first deformation can be implemented by adding UV terms to the Lagrangian. Taken from [Heydeman_2024], the deformation is given by
@@
The Hamiltonian is a sum of multiple Hamiltonians with different numbers of fermions, $q$ and $\tilde{q}_{i}$. If $\tilde{q}_{i}<q$, the second term acts as a relevant deformation controlled by the dimensionless couplings $\epsilon_{i}$ [Anninos_2023, Chapman_2024, Hamdan2025]. This picture is modified slightly when moving to the $\mathcal{N}=2$ model. In this case, a class of relevant deformations can be written
@@
$H=H_{\mathcal{N}=2}^{q=3}+\epsilon\sum_{ij}J_{ij}c_{i}^{\dagger}c_{j}$ (2.21)
@@
The spectral and level spacing densities as a function of the mass deformation strength are shown in Figure 5. Even at small values of the deformation, the BPS degeneracy and energy gap are eliminated. The energy is no longer bounded from below by zero. The level spacing density resembles a Poisson distribution starting at low values of the deformation, and retains its shape for larger values. At higher values of $\epsilon$, the spectral density settles into a Gaussian with positive mean. Figure 4 also shows how the mean gap ratio $\langle r\rangle$ varies with the mass deformation strength $\epsilon$. As soon as the deformation is turned on, the gap ratio drops sharply as the energy gap is eliminated and the spectrum incorporates energy levels from the integrable term. As $\epsilon$ increases further, the integrable term becomes more dominant, and the gap ratio increases to $\langle r\rangle\approx 0.39$, consistent with integrable systems.
@@ 3.1 Krylov Space
Under time evolution by a Hamiltonian, a generic operator will trace out a trajectory in Hilbert space. In many instances, this operator will not explore the entire Hilbert space, but instead will be confined to a subspace. Krylov space is one way of defining this subspace in which the dynamics unfolds [Nandy_2024]. To develop these ideas, consider a Hermitian operator $\mathcal{O}$ evolving in time under the action of a time-independent Hamiltonian $H$. In the Heisenberg picture, the dynamics of the operator is governed by the equation
@@
In addition to the Krylov basis, another output of the Lanczos algorithm is a set of non-negative coefficients $\{b_{n}\}_{n=0}^{D_{K}-1}$, called the Lanczos coefficients, which completely determine the dynamics of the operator in the basis. In many applications, the Lanczos algorithm is numerically unstable due to the accumulation of rounding errors. The effect is that after a few iterations, orthogonality of the Krylov basis may be lost. To ensure orthogonality, techniques like full orthogonalization or partial reorthogonalization can be used [Rabinovici_2021].
@@
Generically, the behavior of the Lanczos coefficients can be divided into three distinct regimes [Nandy_2024]. In the first regime, called the Lanczos ascent, the Lanczos coefficients grow initially until $n\sim O(S)$, where $S$ is the thermodynamic entropy. In the second regime, called the Lanczos plateau, the coefficients plateau at $b_{n}\sim O(\Lambda S)$, where $\Lambda$ is the energy scale. In the final regime, called the Lanczos descent, the $b_{n}$ decrease at exponentially large $n\sim O(e^{S})$.
According to the operator growth hypothesis, in the Lanczos ascent regime, the Lanczos coefficients should grow linearly in $n$ for a generic chaotic system in $d$ dimensions, with an additional logarithmic correction in one-dimensional systems [Parker_2019]. Explicitly, the hypothesis states that for a chaotic system
@@
In general, the Lanczos coefficients need to be computed numerically, but they can be computed exactly for the SYK model when the number of interactions $q$ is large. This can be done using the moments method, starting from the autocorrelation function $C(t)\equiv(\mathcal{O}(t)|\mathcal{O})$ [Viswanath_2013]. To compute $C(t)$ in the large-$q$ limit, an expansion in $1/q$ can be performed:
@@
For $\mathcal{N}=0$ SYK at infinite temperature, $g(t)$ satisfies [Nandy_2024, Parker_2019, Heydeman_2024, Roberts_2018]
@@
From the autocorrelation function, the moments $\mu_{2n}=C^{(2n)}(0)$ can be computed. The Lanczos coefficients can then be calculated from the moments using a recursive algorithm [Viswanath_2013]. The result is [Nandy_2024, Parker_2019]
@@
Similar calculations can be performed for the $\mathcal{N}=2$ SYK model. In this case, the first correction $g(t)$ satisfies [Heydeman_2024]
@@
The leading behavior in $n$ is unmodified, $\alpha=\mathcal{J}$. After the irrelevant deformation of Equation (2.16), the function $g(t)$ satisfies [Heydeman_2024]
@@
This equation is equivalent to that of $\mathcal{N}=0$ SYK with $q$-body interactions deformed by a quadratic term, up to an overall rescaling of the coupling $\mathcal{J}\to\mathcal{J}/2$. Although an analytical solution is not known, the equation can be solved numerically or perturbatively in $\epsilon$ [Chapman_2024, Garcia_Garcia_2018].
@@ 3.2 Krylov Dimension
In Krylov subspace methods, an important role is played by the Krylov dimension $D_{K}$. Formally, $D_{K}$ is the cardinality of the maximal set of linearly independent vectors in the set $\{\mathcal{L}^{n}\mathcal{O}\}_{n=0}^{\infty}$. One can represent these vectors in a convenient matrix form using the basis $|\omega_{ab})\equiv|E_{a}\rangle\langle E_{b}|$ formed by the eigenstates of the Hamiltonian [Rabinovici_2021]. In this basis, the $n$th application of the Liouvillian to the operator $\mathcal{O}$ can be expressed as
@@
It is also possible to understand the Krylov dimension in terms of the number of eigenspaces of the Liouvillian over which the operator $\mathcal{O}$ has non-vanishing projection [Rabinovici_2021]. Suppose that, for some set of indices $I$, the eigenvalue of the Liouvillian is degenerate:
@@ 3.3 Krylov Complexity and Entropy
Like Krylov complexity, Krylov entropy increases during the ascent phase. In general, beyond the scrambling time, Krylov complexity and entropy are related logarithmically: $S(t)\sim\log K(t)$ [Barbon_2019].
@@
For the SYK model, Krylov complexity can be computed analytically in the large-$q$ limit. In $\mathcal{N}=0$, the Lanczos coefficients of Equation (3.11) can be substituted into Equation (3.6) to produce the Krylov-basis wavefunctions. Substituting the result into Equation (3.21), Krylov complexity is given by [Nandy_2024]
@@ 4 Results
This section presents numerical results of Krylov complexity for the $\mathcal{N}=2$ SYK model and its deformations. In Section 4.1, results for the energy degeneracy and Krylov dimension are presented. Then Section 4.2 analyzes the Lanczos sequence for various initial operators. Section 4.3 concludes with results for Krylov complexity and entropy. For the most part, the results are presented for a particular choice of system size $N$, but results across other values of $N$ can be found in Appendix A. The code used to generate the results in this section and the appendices is available on the arXiv page of this paper.
@@ 4.2 Lanczos Sequence
It is also possible to continue the Lanczos sequences to larger values of $n$, up to the eventual termination of the sequences. To accomplish this, the partial reorthogonalization version of the Lanczos algorithm was used. Appendix B reviews the differences between the full orthogonalization and partial reorthogonalization algorithms, and compares their performance on $\mathcal{N}=2$ SYK. Figure 9 shows the full Lanczos sequences of the hopping operator for $\mathcal{N}=2$ SYK with UV and mass deformations. The system size is $N=8$. Without deformation, the sequence terminates after $n\sim 8{,}600$ iterations. Deforming the model extends the lifetime of the Lanczos sequence. Under the irrelevant deformation, the sequence terminates after $n\sim 3.2\times 10^{4}$ iterations, across all values of the deformation strength. With the mass deformation, the sequence terminates after $n\sim 1.3\times 10^{4}$ iterations, also irrespective of the deformation strength. The duration of the plateau is also significantly smaller under the mass deformation.
Figure 9: The complete sequence of Lanczos coefficients $b_{n}$ of $\mathcal{N}=2$ SYK for the hopping operator. Each curve corresponds to the average of ten realizations of the Hamiltonian, with shaded regions indicating the standard deviation. The values $N=8$, $q=3$, and $J=1$ were used throughout.
The Lanczos sequences for $\mathcal{N}=2$ SYK exhibit several interesting features absent in $\mathcal{N}=0$ SYK. Two of these features are shown in Figure 10. In the first feature, the Lanczos coefficients oscillate between two branches, before collapsing to a single branch [Avdoshkin2019]. In the second feature, the Lanczos coefficients collapse, then split back into two branches. These features are present in single realizations of the Hamiltonian, and are not artifacts of averaging over multiple realizations. One possible explanation for these features is that they arise from the operator getting trapped in a small subspace of the overall Hilbert space. In this subspace, the operator might fail to develop large orthogonal directions, reducing the Lanczos coefficients. Once the operator escapes the subspace, the Lanczos coefficients are free to grow again.
(a) $N=8$
(b) $N=9$
Figure 10: Two patterns in the Lanczos coefficients of $\mathcal{N}=2$ SYK for the hopping operator: the “collapse” and “the bridge.” In the collapse, oscillations between two branches of Lanczos coefficients collapse to a single branch. In the bridge, the coefficients collapse to a single branch over a short range before splitting back into two branches. Both patterns were observed for single realizations of the Hamiltonian under the UV deformation with $\epsilon=0.6$.
@@ 4.3 Krylov Complexity and Entropy
Once the Lanczos coefficients have been computed, the dynamics of the operator in the Krylov lattice are fully determined. Given the Lanczos sequence $b_{n}$, the Krylov basis wavefunctions $\varphi_{n}(t)$ were computed by integrating Equation (3.6) numerically out to $t=10$ s, with a spacing of $\Delta t=5\times 10^{-4}$ s in batches of $5\times 10^{-3}$ s. The results were substituted into Equation (3.21) to compute the Krylov complexity $K(t)$, which corresponds to the average position of the operator in the Krylov lattice. This procedure was repeated for each realization of the Hamiltonian, and the results were averaged together at the very end. This quenched disorder averaging is typical in studies of Krylov complexity. If the Lanczos sequences were averaged before computing Krylov complexity, the wavepacket may not randomize efficiently before reaching the end of the Krylov lattice, which could introduce large unphysical oscillations in the complexity [Rabinovici_2021]. In the scenarios considered here, the results were not sensitive to the order of the averaging.
@@
Figure 11 shows Krylov complexity for the different initial operators as the deformation strength is varied. The shaded area around each curve represents the standard deviation from the quenched disorder averaging. For both types of deformations, the complexity grows initially quadratically, then starts growing linearly. In the notation of Table 1, this corresponds to the Krylov exponent $\delta$ transitioning from a half to zero. The transition was observed by computing the Krylov exponent over time based on smoothed estimates of Krylov complexity, and its first and second derivatives. Appendix C discusses the estimation in more detail. The quadratic growth of Krylov complexity is a departure from exponential growth expected from chaotic systems. Initial quadratic growth has also been observed in systems with different symmetries, including SU(2) and Heisenberg-Weyl symmetry [Caputa_2021, Guo_2022]. Quadratic growth is also a generic consequence of working at finite $N$, leading to a finite-dimensional Hilbert space that can stifle the growth of the Lanczos sequence and Krylov complexity. This behavior is also observed in non-supersymmetric Majorana SYK.
@@
Figure 13: The complete profile of Krylov complexity $K(t)$ of $\mathcal{N}=2$ SYK for the hopping operator. The effects of UV and mass deformations on complexity are shown as the deformation strength $\epsilon$ is varied. The values $N=8$, $q=3$, and $J=1$ were used throughout.
Table 2: Comparison of Heisenberg time $t_{H}$, saturation Krylov complexity $K_{\text{sat}}$, and saturation entropy $S_{\text{sat}}$ for different deformation strengths $\epsilon$. Results are shown separately for UV and mass deformations. Also shown are the results for $\mathcal{N}=0$ SYK.
@@
$\epsilon$ $t_{H}$ $K_{\text{sat}}$ $K_{\text{sat}}/D_{K}^{\text{max}}$ $\log K_{\text{sat}}$ $S_{\text{sat}}$ $t_{H}$ $K_{\text{sat}}$ $K_{\text{sat}}/D_{K}^{\text{max}}$ $\log K_{\text{sat}}$ $S_{\text{sat}}$
0.0 988 315 0.22 5.75 5.47 988 315 0.22 5.75 5.47
0.2 23,736 12,552 0.77 9.44 9.35 13,539 5,473 0.08 8.61 8.68
0.4 14,967 11,733 0.72 9.37 9.50 10,644 5,404 0.08 8.59 8.67
0.6 6,071 10,691 0.66 9.28 9.55 8,537 5,396 0.08 8.59 8.67
$\mathcal{N}=0$ 7,569 5,464 0.09 8.61 8.66 7,569 5,464 0.09 8.61 8.66
Krylov complexity can also be computed from the full sequence of Lanczos coefficients, which allows the behavior to be studied at late times. In this case, the hopping equation was integrated out to $t=5\times 10^{5}$ s with $\Delta t=10^{-3}$ s in batches of $500$ s. Figure 13 shows the full profile of Krylov complexity for the hopping operator under both UV and mass deformations for $N=8$, plotted on a log-log scale. In this scale, the transition between quadratic and linear growth corresponds to a decrease in slope from two to one, which is discernible as an elbow in the curves. After the period of quadratic and linear growth, Krylov complexity saturates to a constant value $K_{\text{sat}}$ at the Heisenberg time $t_{H}$. The undeformed model saturates at a value of 315 after 988 s. Turning on the irrelevant deformation delays the Heisenberg time, while increasing the saturation value significantly. The effect of the mass deformation is similar, but the changes are less pronounced given the range of $\epsilon$ considered. For a given value of $\epsilon$, $K_{\text{sat}}$ under the UV deformation is consistently larger than under the mass deformation.
Table 2 shows the Heisenberg time and saturation complexity for different deformation strengths $\epsilon$. The saturation complexity was computed by taking the average of the last 5% of complexity values, while the Heisenberg time was computed as the time when complexity reached 95% of its saturation value. Both of these quantities are sensitive to noise and the threshold values chosen. The table also reports the saturation complexity as a fraction of the maximum Krylov dimension $D_{K}^{\text{max}}$. The undeformed model has a $D_{K}^{\text{max}}$ of 1,4071{,}407, leading to a saturation fraction of $0.22$. The irrelevant deformation increases $D_{K}^{\text{max}}$ to 16,257, which leads to saturation fractions between $0.66$ and $0.77$. The mass deformation increases $D_{K}^{\text{max}}$ to 65,281, leading to a saturation fraction of $0.08$ across all values of $\epsilon$ considered. Although the mass deformation increases the saturation complexity compared to the undeformed model, the saturation complexity is a much smaller fraction of the maximum Krylov dimension. The irrelevant deformation increases both the saturation complexity and the saturation fraction significantly.
@@
The full profile of Krylov entropy is shown in Figure 15. Generally, Krylov entropy grows, plateaus briefly, then continues growing before saturating at late times. Between the plateau and saturation, the entropy grows approximately logarithmically with time, which corresponds to a linear segment in the plots. The logarithmic growth agrees with the theoretical prediction of Table 1, but the entropy cannot be used to fit precise values of the growth rate $\alpha$. As with Krylov complexity, the deformations increase the saturation entropy, while increasing the deformation strength reduces the saturation value. The values of saturation entropy are also reported in Table 2. Comparison with the saturation complexity shows that the saturation entropy is logarithmically related to the saturation complexity, $S_{\text{sat}}\sim\log K_{\text{sat}}$, consistent with the wavefunction spreading evenly across the Krylov basis at late times.
Figure 15: The complete profile of Krylov entropy $S(t)$ of $\mathcal{N}=2$ SYK for the hopping operator. The effects of UV and mass deformations on entropy are shown as the deformation strength $\epsilon$ is varied. The values $N=8$, $q=3$, and $J=1$ were used throughout.

### v2

@@ 2.3 Deformations
The first deformation can be implemented by adding UV terms to the Lagrangian. Taken from [37], the deformation is given by
@@
The Hamiltonian is a sum of multiple Hamiltonians with different numbers of fermions, $q$ and $\tilde{q}_{i}$. If $\tilde{q}_{i}<q$, the second term acts as a relevant deformation controlled by the dimensionless couplings $\epsilon_{i}$ [8, 20, 31]. This picture is modified slightly when moving to the $\mathcal{N}=2$ model. In this case, a class of relevant deformations can be written
@@
$H=H_{\mathcal{N}=2}^{q=3}+\frac{1}{2}\epsilon\sum_{ij}J_{ij}c_{i}^{\dagger}c_{j}$ (2.21)
@@
The spectral and level spacing densities as a function of the mass deformation strength are shown in Figure 5. Even at small values of the deformation, the BPS degeneracy and energy gap are eliminated. The energy is no longer bounded from below by zero. The level spacing density resembles a Poisson distribution starting at low values of the deformation, and retains its shape for larger values. At higher values of $\epsilon$, the spectral density settles into a Gaussian distribution. Figure 4 also shows how the mean gap ratio $\langle r\rangle$ varies with the mass deformation strength $\epsilon$. As soon as the deformation is turned on, the gap ratio drops sharply as the energy gap is eliminated and the spectrum incorporates energy levels from the integrable term. As $\epsilon$ increases further, the integrable term becomes more dominant, and the gap ratio increases to $\langle r\rangle\approx 0.39$, consistent with integrable systems.
@@ 3.1 Krylov Space
Under time evolution by a Hamiltonian, a generic operator will trace out a trajectory in Hilbert space. In many instances, this operator will not explore the entire Hilbert space, but instead will be confined to a subspace. Krylov space is one way of defining this subspace in which the dynamics unfolds [50]. To develop these ideas, consider a Hermitian operator $\mathcal{O}$ evolving in time under the action of a time-independent Hamiltonian $H$. In the Heisenberg picture, the dynamics of the operator is governed by the equation
@@
In addition to the Krylov basis, another output of the Lanczos algorithm is a set of non-negative coefficients $\{b_{n}\}_{n=0}^{D_{K}-1}$, called the Lanczos coefficients, which completely determine the dynamics of the operator in the basis. In many applications, the Lanczos algorithm is numerically unstable due to the accumulation of rounding errors. The effect is that after a few iterations, orthogonality of the Krylov basis may be lost. To ensure orthogonality, techniques like full orthogonalization or partial reorthogonalization can be used [56].
@@
Generically, the behavior of the Lanczos coefficients can be divided into three distinct regimes [50]. In the first regime, called the Lanczos ascent, the Lanczos coefficients grow initially until $n\sim O(S)$, where $S$ is the thermodynamic entropy. In the second regime, called the Lanczos plateau, the coefficients plateau at $b_{n}\sim O(\Lambda S)$, where $\Lambda$ is the energy scale. In the final regime, called the Lanczos descent, the $b_{n}$ decrease at exponentially large $n\sim O(e^{S})$.
According to the operator growth hypothesis, in the Lanczos ascent regime, the Lanczos coefficients should grow linearly in $n$ for a generic chaotic system in $d$ dimensions, with an additional logarithmic correction in one-dimensional systems [51]. Explicitly, the hypothesis states that for a chaotic system
@@
In general, the Lanczos coefficients need to be computed numerically, but they can be computed exactly for the SYK model when the number of interactions $q$ is large. This can be done using the moments method, starting from the autocorrelation function $C(t)\equiv(\mathcal{O}(t)|\mathcal{O})$ [68]. To compute $C(t)$ in the large-$q$ limit, an expansion in $1/q$ can be performed:
@@
For $\mathcal{N}=0$ SYK at infinite temperature, $g(t)$ satisfies [50, 51, 37, 61]
@@
From the autocorrelation function, the moments $\mu_{2n}=C^{(2n)}(0)$ can be computed. The Lanczos coefficients can then be calculated from the moments using a recursive algorithm [68]. The result is [50, 51]
@@
Similar calculations can be performed for the $\mathcal{N}=2$ SYK model. In this case, the first correction $g(t)$ satisfies [37]
@@
The leading behavior in $n$ is unmodified, $\alpha=\mathcal{J}$. After the irrelevant deformation of Equation (2.16), the function $g(t)$ satisfies [37]
@@
This equation is equivalent to that of $\mathcal{N}=0$ SYK with $q$-body interactions deformed by a quadratic term, up to an overall rescaling of the coupling $\mathcal{J}\to\mathcal{J}/2$. Although an analytical solution is not known, the equation can be solved numerically or perturbatively in $\epsilon$ [20, 27].
@@ 3.2 Krylov Dimension
In Krylov subspace methods, an important role is played by the Krylov dimension $D_{K}$. Formally, $D_{K}$ is the cardinality of the maximal set of linearly independent vectors in the set $\{\mathcal{L}^{n}\mathcal{O}\}_{n=0}^{\infty}$. One can represent these vectors in a convenient matrix form using the basis $|\omega_{ab})\equiv|E_{a}\rangle\langle E_{b}|$ formed by the eigenstates of the Hamiltonian [56]. In this basis, the $n$th application of the Liouvillian to the operator $\mathcal{O}$ can be expressed as
@@
It is also possible to understand the Krylov dimension in terms of the number of eigenspaces of the Liouvillian over which the operator $\mathcal{O}$ has non-vanishing projection [56]. Suppose that, for some set of indices $I$, the eigenvalue of the Liouvillian is degenerate:
@@ 3.3 Krylov Complexity and Entropy
Like Krylov complexity, Krylov entropy increases during the ascent phase. In general, beyond the scrambling time, Krylov complexity and entropy are related logarithmically: $S(t)\sim\log K(t)$ [12].
@@
For the SYK model, Krylov complexity can be computed analytically in the large-$q$ limit. In $\mathcal{N}=0$, the Lanczos coefficients of Equation (3.11) can be substituted into Equation (3.6) to produce the Krylov-basis wavefunctions. Substituting the result into Equation (3.21), Krylov complexity is given by [50]
@@ 4 Results
This section presents numerical results of Krylov complexity for the $\mathcal{N}=2$ SYK model and its deformations. In Section 4.1, results for the energy degeneracy and Krylov dimension are presented. Then Section 4.2 analyzes the Lanczos sequence for various initial operators. Section 4.3 concludes with results for Krylov complexity and entropy. For the most part, the results are presented for a particular choice of system size $N$, but results across other values of $N$ can be found in Appendix A. The code used to generate the results in this section and the appendices is available at https://arxiv.org/abs/2511.20769.
@@ 4.2 Lanczos Sequence
It is also possible to continue the Lanczos sequences to larger values of $n$, up to the eventual termination of the sequences. To accomplish this, the partial reorthogonalization version of the Lanczos algorithm was used. Appendix B reviews the differences between the full orthogonalization and partial reorthogonalization algorithms, and compares their performance on $\mathcal{N}=2$ SYK. Figure 9 shows the full Lanczos sequences of the hopping operator for $\mathcal{N}=2$ SYK with UV and mass deformations. The system size is $N=6$. Without deformation, the sequence terminates after $n\sim 33$ iterations, which is below the Krylov bound $D_{K}^{\text{max}}=57$. Deforming the model extends the lifetime of the Lanczos sequence. Under the irrelevant deformation, the sequence terminates after $n\sim 239$ iterations, across all values of the deformation strength, pushing right up to the bound at $D_{K}^{\text{max}}=241$. With the mass deformation, the sequence terminates after $n\sim 859$ iterations, also irrespective of the deformation strength. This is considerably under the bound of $D_{K}^{\text{max}}=4{,}033$.
Figure 9: The complete sequence of Lanczos coefficients $b_{n}$ of $\mathcal{N}=2$ SYK for the hopping operator. Each curve corresponds to the average of ten realizations of the Hamiltonian, with shaded regions indicating the standard deviation. The first value of $\epsilon$ in the legend corresponds to the UV deformation, and the second value to the mass deformation. The values $N=6$, $q=3$, and $J=1$ were used throughout.
The Lanczos sequences for $\mathcal{N}=2$ SYK exhibit interesting features absent in $\mathcal{N}=0$ SYK. An example is shown in Figure 10. In this example, the Lanczos coefficients oscillate between two branches, before collapsing to a single branch [9]. This feature is present in single realizations of the Hamiltonian, and is not an artifact of averaging over multiple realizations. One possible explanation for this feature is that it arises from the operator getting trapped in a small subspace of the overall Hilbert space. In this subspace, the operator might fail to develop large orthogonal directions, reducing the Lanczos coefficients. Once the operator escapes the subspace, the Lanczos coefficients are free to grow again.
Figure 10: An example collapse of the Lanczos coefficients of $\mathcal{N}=2$ SYK for the hopping operator, in which oscillations between two branches collapse to a single branch. This pattern was observed for single realizations of the Hamiltonian under the UV deformation for $N=6$ with $\epsilon=0.6$.
@@ 4.3 Krylov Complexity and Entropy
Once the Lanczos coefficients have been computed, the dynamics of the operator in the Krylov lattice are fully determined. Given the Lanczos sequence $b_{n}$, the Krylov basis wavefunctions $\varphi_{n}(t)$ were computed by integrating Equation (3.6) numerically out to $t=10$ s, with a spacing of $\Delta t=5\times 10^{-4}$ s in batches of $5\times 10^{-3}$ s. The results were substituted into Equation (3.21) to compute the Krylov complexity $K(t)$, which corresponds to the average position of the operator in the Krylov lattice. This procedure was repeated for each realization of the Hamiltonian, and the results were averaged together at the very end. This quenched disorder averaging is typical in studies of Krylov complexity. If the Lanczos sequences were averaged before computing Krylov complexity, the wavepacket may not randomize efficiently before reaching the end of the Krylov lattice, which could introduce large unphysical oscillations in the complexity [56]. In the scenarios considered here, the results were not sensitive to the order of the averaging.
@@
Figure 11 shows Krylov complexity for the different initial operators as the deformation strength is varied. The shaded area around each curve represents the standard deviation from the quenched disorder averaging. For both types of deformations, the complexity grows initially quadratically, then starts growing linearly. In the notation of Table 1, this corresponds to the Krylov exponent $\delta$ transitioning from a half to zero. The transition was observed by computing the Krylov exponent over time based on smoothed estimates of Krylov complexity, and its first and second derivatives. Appendix C discusses the estimation in more detail. The quadratic growth of Krylov complexity is a departure from exponential growth expected from chaotic systems. Initial quadratic growth has also been observed in systems with different symmetries, including SU(2) and Heisenberg-Weyl symmetry [18, 30]. Quadratic growth is also a generic consequence of working at finite $N$, leading to a finite-dimensional Hilbert space that can stifle the growth of the Lanczos sequence and Krylov complexity. This behavior is also observed in non-supersymmetric Majorana SYK.
@@
Figure 13: The complete profile of Krylov complexity $K(t)$ of $\mathcal{N}=2$ SYK for the hopping operator. The first value of $\epsilon$ in the legend corresponds to the UV deformation, and the second value to the mass deformation. The values $N=6$, $q=3$, and $J=1$ were used throughout.
Table 2: Comparison of Heisenberg time $t_{H}$, saturation Krylov complexity $K_{\text{sat}}$, saturation complexity fraction $K_{\text{sat}}^{\text{frac}}\equiv K_{\text{sat}}/D_{K}^{\text{max}}$, and saturation entropy $S_{\text{sat}}$ for different deformation strengths $\epsilon$. Results are shown separately for UV and mass deformations. Also shown are the results for $\mathcal{N}=0$ SYK.
@@
$\epsilon$ $t_{H}$ $K_{\text{sat}}$ $K_{\text{sat}}^{\text{frac}}$ $\log K_{\text{sat}}$ $S_{\text{sat}}$ $\epsilon$ $t_{H}$ $K_{\text{sat}}$ $K_{\text{sat}}^{\text{frac}}$ $\log K_{\text{sat}}$ $S_{\text{sat}}$
0.0 18 8.14 0.14 2.10 2.37 0.0 18 8.14 0.14 2.10 2.37
0.2 933 123.81 0.51 4.82 4.58 1.0 918 341.68 0.08 5.83 5.89
0.4 202 120.64 0.50 4.79 4.64 2.0 327 333.67 0.08 5.81 5.83
0.6 130 117.25 0.49 4.76 4.70 3.0 278 342.84 0.08 5.84 5.81
$\mathcal{N}=0$ 745 320.33 0.10 5.77 5.83 $\mathcal{N}=0$ 745 320.33 0.10 5.77 5.83
Krylov complexity can also be computed from the full sequence of Lanczos coefficients, which allows the behavior to be studied at late times. In this case, the hopping equation was integrated out to $t=10^{4}$ s in batches of $0.01$ s. Figure 13 shows the full profile of Krylov complexity for the hopping operator under both UV and mass deformations for $N=6$, plotted on a log-log scale. In this scale, the transition between quadratic and linear growth corresponds to a decrease in slope from two to one, which is discernible as an elbow in the curves. After the period of quadratic and linear growth, Krylov complexity saturates to a constant value $K_{\text{sat}}$ at the Heisenberg time $t_{H}$. The undeformed model saturates at a value of $8.14$ after $18$ s. Turning on the irrelevant deformation delays the Heisenberg time, while increasing the saturation value significantly. At $\epsilon=0.2$, the model saturates at $123.81$ after $933$ s. As $\epsilon$ is increased further, the Heisenberg time decreases, while the saturation complexity decreases slightly. The mass deformation also increases both the Heisenberg time and saturation complexity relative to the undeformed model.
Table 2 shows the Heisenberg time and saturation complexity for different deformation strengths. The saturation complexity was computed by taking the average of the last 5% of complexity values, while the Heisenberg time was computed as the time when complexity reached 95% of its saturation value. Both of these quantities are sensitive to noise and the threshold values chosen. The table also reports the saturation complexity as a fraction of the maximum Krylov dimension, $K_{\text{sat}}^{\text{frac}}\equiv K_{\text{sat}}/D_{K}^{\text{max}}$. The undeformed model has a $D_{K}^{\text{max}}$ of $57$, leading to a saturation fraction of $0.14$. The irrelevant deformation increases $D_{K}^{\text{max}}$ to $241$, which leads to saturation fractions between $0.49$ and $0.51$. The mass deformation increases $D_{K}^{\text{max}}$ to 4,033, leading to a saturation fraction of $0.08$ across all values of $\epsilon$ considered. Although the mass deformation increases the saturation complexity compared to the undeformed model, the saturation complexity is a much smaller fraction of the maximum Krylov dimension. The irrelevant deformation increases both the saturation complexity and the saturation fraction significantly.
@@
The full profile of Krylov entropy is shown in Figure 15. Generally, Krylov entropy grows, plateaus briefly, then continues growing before saturating at late times. Between the plateau and saturation, the entropy grows approximately logarithmically with time, which corresponds to a linear segment in the plots. The logarithmic growth agrees with the theoretical prediction of Table 1, but the entropy cannot be used to fit precise values of the growth rate $\alpha$. As with Krylov complexity, the deformations increase the saturation entropy. The values of saturation entropy are also reported in Table 2. Comparison with the saturation complexity shows that the saturation entropy is logarithmically related to the saturation complexity, $S_{\text{sat}}\sim\log K_{\text{sat}}$, consistent with the wavefunction spreading evenly across the Krylov basis at late times.
Figure 15: The complete profile of Krylov entropy $S(t)$ of $\mathcal{N}=2$ SYK for the hopping operator. The first value of $\epsilon$ in the legend corresponds to the UV deformation, and the second value to the mass deformation. The values $N=6$, $q=3$, and $J=1$ were used throughout.